OpenStudy (kimberly101):

Shannon buys a used car for $7500. The dealer requires a 15% deposit. Calculate the deposit Shannon has to pay.

OpenStudy (anonymous):

is it 15/100 *7500=1125

OpenStudy (unklerhaukus):

15% of $7500 is \[\frac{15}{100}\times$7500=$\frac{15\times75\times100}{100}\]

OpenStudy (unklerhaukus):

\[=$\frac{15\times75\times\cancel{100}}{\cancel{100}}\]\[=$15\times75\]\[=10\times75+5\times75\]\[=750+375\]\[=1125\]